OpenStudy (anonymous):

how to solve 3/7y-1/4=5/y

OpenStudy (m):

need to know what you're solving for i'm assuming for y? 3/7y - 35/7y = 1/4 -32/7y = 1/4 etc..

OpenStudy (anonymous):

yes it was for y sorry i forgot to put that in there

OpenStudy (anonymous):

You should take the common denominator which is 28y and multiply every term by that. 28y(3) 28y(1) 28y(5) ----- - ------ = ------- 7y 4 y Cancel the denominators into the numerators. 4(3) - 7y(1) = 28(5) 12 - 7y = 140 -7y = 128 y = -18 2/7